# SegChange-R1:Augmented Reasoning for Remote Sensing Change Detection via Large Language Models

### v1

@@ Abstract
Remote sensing change detection is widely used in a variety of fields such as urban planning, terrain and geomorphology analysis, and environmental monitoring, mainly by analyzing the significant change differences of features (e.g., building changes) in the same spatial region at different time phases. In this paper, we propose a large language model (LLM) augmented inference approach (SegChange-R1), which enhances the detection capability by integrating textual descriptive information and aims at guiding the model to segment the more interested change regions, thus accelerating the convergence speed. Moreover, we design a spatial transformation module (BEV) based on linear attention, which solves the problem of modal misalignment in change detection by unifying features from different temporal perspectives onto the BEV space. In addition, we construct the first dataset for building change detection from UAV viewpoints (DVCD ), and our experiments on four widely-used change detection datasets show a significant improvement over existing methods. The code and pre-trained models are available in https://github.com/Yu-Zhouz/SegChange-R1.
@@ 1 Introduction
In remote sensing, change detection (CD) refers to the process of analyzing remote sensing images of the same area acquired at different time phases to identify changes in surface features [2, 24]. CD tasks have a wide range of applications, such as urban expansion monitoring  [47, 50], disaster assessment [1, 43], land use and land cover change analysis [13, 40, 54], and military reconnaissance [20, 23]. However, due to the influence of various factors on remote sensing images, CD still faces numerous challenges.
@@
First, images collected at different time points often exhibit variations in lighting and seasonal changes, leading to significant spectral differences in the same objects across different time phases [6, 32]. Second, inconsistent resolution among multi-source remote sensing data can also affect the accuracy of change information extraction [4]. Additionally, factors such as sensor noise and atmospheric interference introduce image noise, further complicating change modeling [2, 46]. Another critical issue is image registration error; even after preprocessing, minor spatial misalignments may lead to incorrect change judgments [4, 39].
@@
In recent years, methods based on convolutional neural networks (such as FC-EF and FC-Siam-diff) have enhanced the feature consistency of dual-temporal images through twin structures, thereby improving detection performance. To further model long-range dependencies, Transformer-based methods (such as BIT [11] and ChangeFormer [3]) introduce self-attention mechanisms, demonstrating superior performance in multi-scale change modeling. ChangeMamba [26] is based on the Mamba architecture and utilizes a state space model to process long sequences of remote sensing data, improving modeling efficiency. However, most methods remain limited to visual feature extraction, lacking semantic-level understanding and reasoning capabilities, which affects the accuracy and convergence speed of change features. Recently, BEV space modeling has been introduced to unify perspective representation, aiming to address this issue. Additionally, how to combine contextual information (such as text descriptions or geographic tags) to enhance the model’s focus on regions of interest remains an area worth exploring [27].
@@
Recently, the development of large language models (LLMs) has brought new opportunities to this field. In this paper, we propose a new method called SegChange-R1, which utilizes LLMs to combine textual description information with remote sensing images, guiding the model to focus more on regions of interest and thereby improving the detection of significant changes in objects between two time phases. As shown in Fig. 2, specifically, we designed a spatial transformation module (BEV) based on linear attention. This module addresses the issue of modal mismatch in change detection by unifying features from different temporal perspectives into the BEV space. Compared to transformers, the linear attention-based architecture enhances feature expression capabilities by modeling global dependencies, demonstrating linear time training capabilities and more efficient modeling of spatial dependencies, enabling rapid model convergence. In summary, the main contributions of our work are as follows:
@@
  • We have developed a novel semantic-guided enhanced reasoning remote sensing change detector (SegChange-R1) based on a large prophecy model, which generates more accurate location masks by integrating textual description information given two images.
@@
  • To this end, we designed a spatial transformation module (BEV) based on linear attention to address the issue of modal mismatch in change detection.
@@
  • We constructed the first drone-view building change detection dataset (DVSC), which contains 13,800 pairs of change images covering building changes in various scenarios from urban to rural areas.
@@ 2 Related Work
Remote sensing change detection, as an important remote sensing application technology, has undergone a revolutionary evolution from traditional methods to deep learning methods. Early change detection methods primarily relied on pixel-level difference analysis, including image difference methods, ratio methods, change vector analysis, and principal component analysis [2, 24, 40]. While these traditional methods are computationally simple, they are often influenced by various factors such as changes in lighting, seasonal variations, sensor noise, and atmospheric conditions, leading to high false alarm rates and making it difficult to accurately identify genuine changes in land cover.
With the rapid development of deep learning technology, change detection methods based on convolutional neural networks (CNNs) have gradually become the mainstream paradigm. Early deep learning methods such as FC-EF (Fully Convolutional Early Fusion) employed an early fusion strategy, concatenating dual-phase images at the input layer [16]. Subsequently, twin network architectures such as FC-Siam-diff (Fully Convolutional Siamese Difference) and FC-Siam-conc were proposed, which perform parallel processing of dual-phase images using feature extractors with shared weights, followed by feature fusion through difference or concatenation operations [15, 35]. These methods ensure feature consistency through parameter sharing, significantly enhancing the accuracy and robustness of change detection.
To further improve detection performance, researchers have begun exploring more complex network architectures and attention mechanisms. STANet [10] introduced a spatio-temporal attention mechanism, enhancing the model’s ability to focus on changing regions through spatial attention and channel attention modules. DT-CDSCN [19] proposed a dual-task constrained deep twin convolutional network, improving change detection performance through auxiliary training with a semantic segmentation task. IFN [51] designed an interactive feature fusion network, achieving more precise change modeling through multi-level feature interaction. In recent years, attention-based methods such as SNUNet [18] and FCCDN [8] have further advanced this field, enhancing feature representation and fusion capabilities through the design of specialized attention modules.
@@
To better model long-range dependencies and global contextual information, change detection methods based on Transformers have emerged, bringing new breakthroughs to the field. BIT (Binary Change Detection with Transformers) [11] was the first to introduce the Transformer architecture into change detection tasks, using self-attention mechanisms to capture global contextual information, achieving significant performance improvements across multiple benchmark datasets. ChangeFormer [3] further improved the Transformer structure by designing specialized change-aware attention modules and hierarchical feature fusion strategies. SwinSUNet [44] combines the hierarchical feature representation capabilities of the Swin Transformer with a sliding window mechanism, demonstrating outstanding performance when handling multi-scale changes.
While the traditional Transformer architecture excels at modeling global dependencies, its $O(n^{2})$ computational complexity becomes a bottleneck when processing high-resolution remote sensing images. To address this issue, researchers have begun exploring more efficient attention mechanisms and alternative architectures. ChangeMamba [26] designs a change detection architecture with linear complexity based on a state space model (SSM), achieving efficient long sequence modeling through Mamba’s selective scanning mechanism. Performer [12] achieves linear-complexity attention computation through random feature mapping, while Linformer [45] reduces the dimension of the attention matrix via low-rank decomposition. These methods significantly reduce computational costs while maintaining similar performance, making them particularly suitable for processing high-resolution remote sensing image data.
@@
Existing change detection methods have placed greater emphasis on visual representation learning while neglecting the potential of multimodal data. CLIP (Contrastive Language-Image Pre-training) [38] successfully demonstrated the effectiveness of visual-language pre-training across various visual tasks, laying the foundation for multimodal applications in the field of remote sensing. Subsequent works such as FLAVA [41] and ALIGN [22] further explored the possibilities of large-scale multimodal pre-training. Inspired by this, some researchers began to incorporate language information into remote sensing change detection tasks, exploring how text descriptions can guide models to focus on specific types of changes [27, 30, 51, 33]. Recently, the rapid development of large language models (LLMs) has brought new opportunities to remote sensing change detection. Multimodal large language models such as GPT-4V [36], LLaVA [31], and InstructBLIP [14] have demonstrated strong visual understanding and reasoning capabilities. In the field of remote sensing, LLMs are increasingly being applied to tasks such as image description generation, scene understanding, and object detection [7, 48, 42]. In particular, LLMs’ capabilities in spatial reasoning and regional localization offer new possibilities for remote sensing change detection. By combining visual features with natural language descriptions, LLMs can better understand the semantic meaning of changes, thereby guiding models to focus on regions of interest to users.
However, the application of existing multimodal fusion methods in remote sensing change detection remains relatively limited. Most methods still rely on comparisons of pure visual features, lacking a deep understanding of semantic-level changes. How to effectively integrate the reasoning capabilities of LLMs into the change detection process and achieve a deep fusion of visual features and language instructions remains a critical issue that needs to be addressed.
@@
Spatial alignment and efficient architecture design are two key technical challenges in remote sensing change detection. Due to various interference factors during remote sensing image acquisition, such as sensor position differences, changes in shooting angles, and atmospheric conditions, images from different time phases often exhibit minor but non-negligible spatial misalignment issues [5, 39]. Such misalignment can lead to erroneous change detection results even at the sub-pixel level, particularly in edge regions and the detection of small targets.
Bird’s Eye View (BEV) representation, as a unified spatial representation method, has been widely applied in autonomous driving and 3D object detection fields [17, 37, 53]. The core advantage of BEV representation lies in its ability to unify data from different perspectives and sensors into a single spatial coordinate system, effectively addressing perspective changes and spatial misalignment issues. LSS (Lift, Splat, Shoot) [37] converts perspective view features into BEV representation through depth estimation, while BEVFormer [29] further incorporates temporal information modeling. Recently, some researchers have begun to incorporate BEV representations into remote sensing change detection tasks to address registration errors between multi-temporal images [28, 9]. By mapping image features from different temporal phases into a unified BEV space, the impact of spatial misalignment on change detection can be reduced.
Figure 1: An overview of the SegChange-R1 Architecture. Given two images before and after a change, they are passed through a shared BackBone to extract multi-scale features, which are then input into the difference module for feature difference extraction. The features are then sent to the FPN layer for fusion. Finally, the resulting feature maps and text semantic features are input into the Mask Head to predict the feature map of the changed area.
@@ 3 Proposed Method
In this work, we propose SegChange-R1, a simple yet powerful geospatial pixel inference baseline. As shown in Fig. 1, our architecture consists of three main components: a pre-trained text-image encoder, a BEV space transformer, and a masked Decoder.
@@ 3.1 Encoder
Vision Encoder.
Remote sensing images exhibit significant scale variation characteristics, ranging from sub-meter-scale small objects to kilometer-scale large-scale geographical structures, posing a significant challenge to the model’s multi-scale modeling capabilities [53]. Additionally, the dense distribution of small targets in high-resolution remote sensing images requires the model to retain as much spatial detail information as possible during feature extraction [12]. However, current mainstream Vision Transformer-based encoders (such as CLIP [3] and SAM [18, 41]) have limitations when handling such tasks: their fixed window mechanism and aggressive downsampling strategy easily lead to the loss of small-scale target information, limiting the model’s perception capabilities in complex remote sensing scenarios. To address these issues, we adopted the Swin Transformer backbone network [1], which employs a sliding window mechanism to model local attention, thereby enhancing the model’s ability to capture fine-grained features while maintaining computational efficiency. Based on this, we constructed a progressive multi-scale feature extraction framework, generating feature maps at resolutions of 1/4, 1/8, 1/16 and 1/32 of the original input image, denoted as $v_{h}\in[1,4]$, thereby balancing spatial resolution and semantic abstraction at different levels. Additionally, our implementation supports multiple backbone network options (including ResNet50 [21], Swin-Transformer [34], and HGNetv2 [52]) to accommodate diverse computational resource and accuracy requirements across various application scenarios.
Text Encoder.
Text prompts play a crucial role in remote sensing change detection tasks, as they provide models with semantic information about the type of changes in target objects, thereby guiding the model to focus on changes in specific features. Recent advancements in vision-language models have demonstrated that incorporating text semantic information to enhance visual understanding capabilities is highly effective [38]. Notably, ChangeCLIP [51] leverages comprehensive textual semantic information from remote sensing images to enhance visual models’ perception of remote sensing changes, achieving unprecedented performance across multiple benchmark datasets. Additionally, studies such as PromptCC [33] have validated the effectiveness of using large language models (LLMs) for text prompt encoding, where multi-prompt learning strategies significantly improve the accuracy of language generation.
To effectively bridge the semantic gap between text descriptions and visual features, we introduce a text encoder based on a pre-trained LLM, Microsoft/Phi-1.5. This encoder converts input text descriptions into semantically rich embedding vectors, which are then integrated with the visual encoder through a deep feature fusion mechanism. This design enables the model to focus on specific types of land cover changes guided by semantic constraints derived from natural language descriptions. Additionally, we have implemented a dynamic sequence length control mechanism, enabling text embeddings to adapt to the requirements of different downstream tasks.
@@ 3.2 BEV Space Converter
In remote sensing change detection, modal mismatch is a major challenge. To address this issue, we propose a novel BEV space transformation module that plays a key role in our SegChange-R1 framework. This module is designed to address the inherent challenges of modal mismatch when processing remote sensing data from different time periods. The module is based on a linear attention mechanism, enabling efficient and effective feature transformation. The idea is to transform features from different temporal perspectives into a unified BEV (bird’s-eye view) space, thereby enabling more effective comparison and analysis of changes.
The BEV Space Converter takes features from multiple time phases as input. These features are first projected into a latent space using linear transformations. Mathematically, this can be represented as follows:
@@
Using these attention weights, the features are aggregated to form a unified representation in the BEV space. This allows the model to effectively address the modality misalignment problem and better capture the changes between different time phases.
The BEV Space Converter not only enhances the model’s ability to detect changes but also provides a more robust and interpretable feature representation for remote sensing change detection tasks.
@@
Table 1: Dataset Information.Three publicly available building change detection benchmark datasets and details of the dataset we propose.

### v2

# SegChange-R1: LLM-Augmented Remote Sensing Change Detection

@@ Abstract
Remote sensing change detection is used in urban planning, terrain analysis, and environmental monitoring by analyzing feature changes in the same area over time. In this paper, we propose a large language model (LLM) augmented inference approach (SegChange-R1), which enhances the detection capability by integrating textual descriptive information and guides the model to focus on relevant change regions, accelerating convergence. We designed a linear attention-based spatial transformation module (BEV) to address modal misalignment by unifying features from different times into a BEV space. Furthermore, we introduce DVCD, a novel dataset for building change detection from UAV viewpoints. Experiments on four widely-used datasets demonstrate significant improvements over existing method The code and pre-trained models are available in https://github.com/Yu-Zhouz/SegChange-R1.
@@ 1 Introduction
Change detection (CD) in remote sensing analyzes images of the same area acquired at different times to identify surface feature changes [2, 24]. CD tasks are used for urban expansion monitoring [47, 50], disaster assessment [1, 43], land use and land cover change analysis [13, 40, 54], and military reconnaissance [20, 23]. However, CD faces challenges due to various factors affecting remote sensing images.
@@
First, images from different time points often exhibit lighting and seasonal variations, causing spectral differences in the same objects [6, 32]. Second, inconsistent resolution in multi-source data can affect the accuracy of change extraction [4]. Sensor noise and atmospheric interference also introduce image noise, complicating change modeling [2, 46]. Image registration errors are another critical issue; even after preprocessing, minor misalignments can lead to incorrect change judgments [4, 39].
@@
Recently, convolutional neural networks (CNNs) such as Fully Convolutional Early Fusion (FC-EF) and FC-Siam-diff have enhanced feature consistency in dual-temporal images using twin structures, improving detection performance. Transformer-based methods, such as BIT [11] and ChangeFormer [3], use self-attention to model long-range dependencies, improving multi-scale change modeling. ChangeMamba [26] is based on the Mamba architecture and utilizes a state space model to process long sequences of remote sensing data, improving modeling efficiency. However, most methods are limited to visual feature extraction and lack semantic understanding, affecting the accuracy and convergence speed of change features. Recently, Bird’s Eye View (BEV) space modeling has been introduced to unify perspective representation, aiming to address this issue. Combining contextual information, such as text descriptions or geographic tags, to focus the model on regions of interest remains an area for exploration [27].
@@
Recently, the rapid development of LLMs has brought new opportunities to this field. We propose SegChange-R1, which uses LLMs to combine textual descriptions with remote sensing images. This guides the model to focus on regions of interest, improving the detection of significant changes between time phases. As shown in Fig. 2, we designed a spatial transformation module (BEV) based on linear attention. This module unifies features from different temporal perspectives into a shared BEV space, addressing modal mismatch in change detection. Compared to transformers, the linear attention-based architecture enhances feature expression by modeling global dependencies, enabling linear-time training, efficient spatial dependency modeling, and rapid convergence. In summary, the main contributions of our work are as follows:
@@
  • We developed SegChange-R1, a novel semantic-guided remote sensing change detector that uses a large language model to generate accurate location masks by integrating textual descriptions of two images.
@@
  • We designed a linear attention-based spatial transformation module (BEV) to address modal mismatch in change detection.
@@
  • We introduce DVSC, a drone-view building change detection dataset containing 13,800 image pairs of building changes in diverse urban and rural scenarios.
@@ 2 Related Work
Remote sensing change detection has evolved from traditional methods to deep learning methods. Early change detection methods primarily relied on pixel-level difference analysis, including image difference methods, ratio methods, change vector analysis, and principal component analysis [2, 24, 40]. While these traditional methods are computationally simple, they are often influenced by various factors such as changes in lighting, seasonal variations, sensor noise, and atmospheric conditions, leading to high false alarm rates and making it difficult to accurately identify genuine changes in land cover.
With the development of deep learning, convolutional neural networks (CNNs) have become the mainstream paradigm for change detection. Early deep learning methods such as FC-EF (Fully Convolutional Early Fusion) employed an early fusion strategy, concatenating dual-phase images at the input layer [16]. Subsequently, twin network architectures such as FC-Siam-diff (Fully Convolutional Siamese Difference) and FC-Siam-conc were proposed, which perform parallel processing of dual-phase images using feature extractors with shared weights, followed by feature fusion through difference or concatenation operations [15, 35]. These methods ensure feature consistency through parameter sharing, enhancing the accuracy and robustness of change detection.
To improve detection performance, researchers have explored complex network architectures and attention mechanisms. STANet [10] introduced a spatio-temporal attention mechanism, enhancing the model’s ability to focus on changing regions through spatial attention and channel attention modules. DT-CDSCN [19] proposed a dual-task constrained deep twin convolutional network, improving change detection performance through auxiliary training with a semantic segmentation task. IFN [51] designed an interactive feature fusion network, achieving more precise change modeling through multi-level feature interaction. In recent years, attention-based methods such as SNUNet [18] and FCCDN [8] have further advanced this field, enhancing feature representation and fusion capabilities through the design of specialized attention modules.
@@
To better model long-range dependencies and global context, Transformer-based change detection methods have emerged. BIT (Binary Change Detection with Transformers) [11] was the first to introduce the Transformer architecture into change detection tasks, using self-attention mechanisms to capture global contextual information, achieving significant performance improvements across multiple benchmark datasets. ChangeFormer [3] further improved the Transformer structure by designing specialized change-aware attention modules and hierarchical feature fusion strategies. SwinSUNet [44] combines the hierarchical feature representation capabilities of the Swin Transformer with a sliding window mechanism, demonstrating outstanding performance when handling multi-scale changes.
While Transformers excel at modeling global dependencies, their $O(n^{2})$ complexity becomes a bottleneck when processing high-resolution remote sensing images. To address this issue, researchers have begun exploring more efficient attention mechanisms and alternative architectures. ChangeMamba [26] designs a change detection architecture with linear complexity based on a state space model (SSM), achieving efficient long sequence modeling through Mamba’s selective scanning mechanism. Performer [12] achieves linear-complexity attention computation through random feature mapping, while Linformer [45] reduces the dimension of the attention matrix via low-rank decomposition. These methods reduce computational costs while maintaining performance, making them suitable for high-resolution remote sensing images.
@@
Existing change detection methods emphasize visual representation learning, neglecting the potential of multimodal data. CLIP (Contrastive Language-Image Pre-training) [38] successfully demonstrated the effectiveness of visual-language pre-training across various visual tasks, laying the foundation for multimodal applications in the field of remote sensing. Subsequent works such as FLAVA [41] and ALIGN [22] further explored the possibilities of large-scale multimodal pre-training. Inspired by this, some researchers began to incorporate language information into remote sensing change detection tasks, exploring how text descriptions can guide models to focus on specific types of changes [27, 30, 51, 33]. Recently, the rapid development of LLMs has brought new opportunities to remote sensing change detection. Multimodal large language models such as Generative Pre-trained Transformer 4 with Vision (GPT-4V) [36], LLaVA [31], and InstructBLIP [14] have demonstrated strong visual understanding and reasoning capabilities. In the field of remote sensing, LLMs are increasingly being applied to tasks such as image description generation, scene understanding, and object detection [7, 48, 42]. In particular, LLMs’ capabilities in spatial reasoning and regional localization offer new possibilities for remote sensing change detection. By combining visual features with natural language descriptions, LLMs can better understand the semantics of changes, guiding models to focus on regions of interest.
However, multimodal fusion in remote sensing change detection remains limited. Most methods still rely on visual feature comparisons, lacking a deep understanding of semantic changes. Effectively integrating LLM reasoning into change detection and achieving deep fusion of visual features and language remains a critical issue.
@@
Spatial alignment and efficient architecture are key technical challenges in remote sensing change detection. Due to various interference factors during remote sensing image acquisition, such as sensor position differences, changes in shooting angles, and atmospheric conditions, images from different time phases often exhibit minor but non-negligible spatial misalignment issues [5, 39]. This misalignment can lead to erroneous change detection, even at the sub-pixel level, particularly in edge regions and for small targets.
Bird’s Eye View (BEV) representation, as a unified spatial representation method, has been widely applied in autonomous driving and 3D object detection fields [17, 37, 53]. BEV’s advantage lies in unifying data from different perspectives and sensors into a single spatial coordinate system, addressing perspective changes and spatial misalignment. LSS (Lift, Splat, Shoot) [37] converts perspective view features into BEV representation through depth estimation, while BEVFormer [29] further incorporates temporal information modeling. Recently, some researchers have begun to incorporate BEV representations into remote sensing change detection tasks to address registration errors between multi-temporal images [28, 9]. By mapping image features from different temporal phases into a unified BEV space, the impact of spatial misalignment on change detection can be reduced.
Figure 1: An overview of the SegChange-R1 Architecture. Given two images before and after a change, they are passed through a shared BackBone to extract multi-scale features, which are then input into the difference module for feature difference extraction. The features are then sent to the Feature Pyramid Network (FPN) layer for fusion. Finally, the resulting feature maps and text semantic features are input into the Mask Head to predict the feature map of the changed area.
@@ 3 Proposed Method
We propose SegChange-R1, a geospatial pixel inference baseline. As shown in Fig. 1, our architecture includes a pre-trained text-image encoder, a BEV space transformer, and a masked decoder.
@@ 3.1 Encoder
Vision Encoder
Remote sensing images exhibit significant scale variation, from sub-meter objects to kilometer-scale structures, challenging the model’s multi-scale modeling capabilities [53]. Additionally, the dense distribution of small targets in high-resolution images requires the model to retain spatial detail during feature extraction [12]. However, mainstream Vision Transformer encoders (such as CLIP [3] and Segment Anything Model (SAM) [18, 41]) have limitations: their fixed window mechanism and aggressive downsampling can lead to the loss of small-scale target information, limiting perception in complex remote sensing scenarios. To address these issues, we adopted the Swin Transformer [1], which uses a sliding window to model local attention, enhancing the capture of fine-grained features while maintaining efficiency. Based on this, we constructed a progressive multi-scale feature extraction framework, generating feature maps at resolutions of 1/4, 1/8, 1/16 and 1/32 of the original input image, denoted as $v_{h}\in[1,4]$, thereby balancing spatial resolution and semantic abstraction at different levels. Additionally, our implementation supports multiple backbones (including ResNet50 [21], Swin-Transformer [34], and HGNetv2 [52]) to accommodate diverse computational and accuracy requirements.
Text Encoder
Text prompts are crucial in remote sensing change detection, providing semantic information about changes and guiding the model to focus on specific features. Vision-language models have demonstrated that incorporating text semantics effectively enhances visual understanding [38]. ChangeCLIP [51] leverages comprehensive text semantics from remote sensing images to enhance visual models’ perception of changes, achieving strong performance on benchmark datasets. Additionally, studies such as PromptCC [33] validate the effectiveness of using LLMs for text prompt encoding; multi-prompt learning improves language generation accuracy.
To bridge the semantic gap between text and visual features, we introduce a text encoder based on a pre-trained Large Language Model (LLM), Microsoft/Phi-1.5. This encoder converts text descriptions into semantically rich embeddings, integrated with the visual encoder through deep feature fusion. This enables the model to focus on specific land cover changes guided by semantic constraints from text. Additionally, we implemented a dynamic sequence length control, enabling text embeddings to adapt to downstream tasks.
@@ 3.2 BEV Space Converter
Modal mismatch is a major challenge in remote sensing change detection. To address this, we propose a BEV space transformation module in our SegChange-R1 framework. This module addresses modal mismatch when processing data from different times and is based on a linear attention mechanism for efficient feature transformation. The module transforms features from different times into a shared BEV space, enabling effective comparison and change analysis.
The BEV Space Converter takes features from multiple times as input, projecting them into a latent space using linear transformations. Mathematically, this can be represented as follows:
@@
Using these attention weights, the features are aggregated into a unified BEV representation, allowing the model to address modality misalignment and capture changes between different times.
The BEV Space Converter enhances change detection and provides a robust, interpretable feature representation.
@@
Table 1: Datasets used for building change detection. This table summarizes the characteristics of three publicly available benchmark datasets, along with details of the newly proposed dataset.